# SegEarth-OV: Towards Traning-Free Open-Vocabulary Segmentation for Remote Sensing Images

### v1

@@ Abstract
Remote sensing image plays an irreplaceable role in fields such as agriculture, water resources, military, and disaster relief. Pixel-level interpretation is a critical aspect of remote sensing image applications; however, a prevalent limitation remains the need for extensive manual annotation. For this, we try to introduce open-vocabulary semantic segmentation (OVSS) into the remote sensing context. However, due to the sensitivity of remote sensing images to low-resolution features, distorted target shapes and ill-fitting boundaries are exhibited in the prediction mask. To tackle this issue, we propose a simple and general upsampler, SimFeatUp, to restore lost spatial information in deep features in a training-free style. Further, based on the observation of the abnormal response of local patch tokens to [CLS] token in CLIP, we propose to execute a straightforward subtraction operation to alleviate the global bias in patch tokens. Extensive experiments are conducted on 17 remote sensing datasets spanning semantic segmentation, building extraction, road detection, and flood detection tasks. Our method achieves an average of 5.8%, 8.2%, 4%, and 15.3% improvement over state-of-the-art methods on 4 tasks. All codes are released.
@@ 1 Introduction
Remote sensing imagery has changed the way humans observe and understand the Earth. It enables us to monitor land cover/use types, respond effectively to natural disasters (e.g., fires, earthquakes, floods), gain insight into food and water resources, etc. Among the 17 Sustainable Development Goals (SDGs) issued by the United Nations, remote sensing imagery can provide important data support for several goals including “Zero Hunger”, “Clean Water and Sanitation”, “Industry, Innovation and Infrastructure”, “Climate Action”, “Life on Land” [61]. Notably, remote sensing data can be considered as a distinct modality in machine learning. It involves more diverse spatial resolutions (from centimeters to kilometers), temporal dimensions (from hours to decades), and object perspectives (overhead and oriented) than natural images. Therefore, solutions designed for other data modalities (e.g., natural images) may be sub-optimal for remote sensing data [53].
@@
  • We present SimFeatUp, an general upsampler for training-free OVSS which robustly upsamples low-resolution (LR) features and maintains semantic consistency with image content.
  • We present an extremely simple and straightforward way to alleviate the global bias problem of CLIP, i.e., executing subtraction operations of local and global tokens.
@@ 2 Related Work
Open-Vocabulary Semantic Segmentation. As VLMs have shown remarkable zero-shot inference in image classification, which naturally extends to semantic segmentation. They empower the segmentation pipeline to recognize seen and unseen categories, and users can segment almost any category in an image using prompt vocabulary [68, 81]. We divide current CLIP-based OVSS methods into two groups: training-required and training-free. The former allows models to be trained on some base classes in a supervised or weakly supervised manner and is therefore more flexible. Typically, some works [16, 48, 52, 42] try to train a localization-aware CLIP which can naturally make dense predictions, while others [30, 13, 74, 73, 11, 38] select a subset of the CLIP’s pre-trained parameters and/or introduce a limited number of trainable parameters into the frozen CLIP, i.e., fine-tuning the CLIP to adapt to dense prediction on base classes. Still, training-free OVSS methods emphasize tapping into CLIP’s inherent localization capabilities with limited surgery of features or structures. MaskCLIP [79] pioneers the removal of query and key projections at the attention pooling layer of CLIP’s image encoder. Following it, subsequent studies [34, 63, 3, 28] adequately explore self-self attention (i.e., q-q, k-k or v-v self-attention), and these modifications somewhat mitigate noisy activations and spatial invariant perception of CLIP. Another stream [56, 23, 60, 2] is the two-stage method, which first generates category-agnostic mask proposals and then classifies the masks. Besides, some other foundation models (e.g. DINO [6], SAM [26], Stable Diffusion [54]) can be introduced to enhance the localization ability of CLIP, and these explorations also make sense [29, 2, 65].
Different from previous methods, we focus on the inherent characteristics of remote sensing images rather than the general attributes of natural images. Our SimFeatUp component, although it needs to be trained on a few images-only data beforehand, this process is independent of the semantic segmentation process and the trained weights can be used for almost any remote sensing data (like the foundation model in other works [29, 2]), so our method can still be seen as a training-free method.
@@ 3.1 CLIP
$\begin{array}[]{c}\boldsymbol{q}=\operatorname{Emb}_{q}(\boldsymbol{X}),\boldsymbol{k}=\operatorname{Emb}_{k}(\boldsymbol{X}),\boldsymbol{v}=\operatorname{Emb}_{v}(\boldsymbol{X}),\\ \boldsymbol{y}=\boldsymbol{x}+\operatorname{SA}\left(\boldsymbol{q},\boldsymbol{k},\boldsymbol{v}\right),\\ \boldsymbol{z}=\boldsymbol{y}+\operatorname{FFN}(\operatorname{LN}(\boldsymbol{y})),\end{array}$ (1)
@@ 4.1 SimFeatUp
FeatUp provides us with an excellent training paradigm for general upsamplers. However, it lacks some considerations for the traning-free setting, leading it to be sub-optimal for the OVSS task, especially in remote sensing contexts.
@@ 4.2 Alleviating global bias
Figure 5: Comparison before and after alleviating the global bias. (a) is the similarity map of patch tokens and cls tokens, some “non-building” regions also present high response, (b) is the original RGB image. Note that the right-hand histograms stretch the raw values for better presentation.
@@ 5.3 Comparison to State-of-the-art
Since the proposed SegEarth-OV is a traning-free method and there is no previous OVSS method designed for remote sensing images, we select 5 state-of-the-art traning-free OVSS models of natural images for comparison, including vanilla CLIP [51], MaskCLIP [79], SCLIP [63], GEM [3] and ClearCLIP [28].
Semantic segmentation. As listed in Tab. 1, SegEarth-OV achieves the best performance on all 8 semantic segmentation datasets. SegEarth-OV achieves more than 40% mIoU on 5 datasets and more than 50% on the UDD5 dataset, which implies that the OVSS method is usable in remote sensing scenarios. Compared to the previous method, SegEarth-OV achieves a performance gain of more than 5% on 5 datasets and an average gain of 5.8% on 8 datasets. On the iSAID dataset, the mIoU of SegEarth-OV is only 21.7%, which is due to the fine-grained category delineation in this dataset, which covers 16 categories (see Appendix Tab. 7).
@@ 5.4 Ablation Study and Analysis
Remote sensing CLIPs for OVSS. We evaluate the performance of remote sensing CLIPs on OVSS, including RemoteCLIP [36], GeoRSCLIP [77], and SkyCLIP [66], which are trained on 0.8M, 5M, and 5.2M remote sensing data, respectively, without changing the model structure of CLIP. Since these works do not provide the ViT-B/16, we uniformly use ViT-B/32. Hence, we repeat the JBU operation 5 times in SimFeatUp. For fair comparison, we train the respective SimFeatUp for each model. As listed in Tab. 3, RemoteCLIP performs suboptimally to vanilla CLIP, which indicates that a small amount of domain data diminishes the model’s transfer capability. GeoRSCLIP achieves the best performance against SkyCLIP, one possible reason is that the images and descriptions used in GeoRSCLIP are more diverse. Moreover, the OVSS task effectively reflects the model’s discrimination and localization capabilities, and can serve as an evaluation metric for remote sensing VLMs.

### v2

# SegEarth-OV: Towards Training-Free Open-Vocabulary Segmentation for Remote Sensing Images

@@ Abstract
Remote sensing image plays an irreplaceable role in fields such as agriculture, water resources, military, and disaster relief. Pixel-level interpretation is a critical aspect of remote sensing image applications; however, a prevalent limitation remains the need for extensive manual annotation. For this, we try to introduce open-vocabulary semantic segmentation (OVSS) into the remote sensing context. However, due to the sensitivity of remote sensing images to low-resolution features, distorted target shapes and ill-fitting boundaries are exhibited in the prediction mask. To tackle this issue, we propose a simple and general upsampler, SimFeatUp, to restore lost spatial information in deep features in a training-free style. Further, based on the observation of the abnormal response of local patch tokens to [CLS] token in CLIP, we propose to execute a straightforward subtraction operation to alleviate the global bias in patch tokens. Extensive experiments are conducted on 17 remote sensing datasets spanning semantic segmentation, building extraction, road detection, and flood detection tasks. Our method achieves an average of 5.8%, 8.2%, 4.0%, and 15.3% improvement over state-of-the-art methods on 4 tasks. All codes are released.
@@ 1 Introduction
Remote sensing image has changed the way humans observe and understand the Earth. It enables us to monitor land cover/use types, respond effectively to natural disasters (e.g., fires, earthquakes, floods), gain insight into food and water resources, etc. Among the 17 Sustainable Development Goals (SDGs) [^1] https://sdgs.un.org/ issued by the United Nations, remote sensing image can provide important data support for several goals including “Zero Hunger”, “Clean Water and Sanitation”, “Industry, Innovation and Infrastructure”, “Climate Action”, “Life on Land”, etc. [61] Notably, remote sensing data can be considered as a distinct modality in machine learning. It involves more diverse spatial resolutions (from centimeters to kilometers), temporal dimensions (from hours to decades), and object perspectives (overhead and oriented) than natural images. Therefore, solutions designed for other data modalities (e.g., natural images) may be sub-optimal for remote sensing data [53].
@@
  • We propose SimFeatUp, an general feature upsampler for training-free OVSS, which robustly upsamples low-resolution (LR) features and maintains semantic consistency with image content.
  • We propose an extremely simple and straightforward way to alleviate the global bias problem of CLIP, i.e., executing subtraction operations of local and global tokens.
@@ 2 Related Work
Open-Vocabulary Semantic Segmentation. As VLMs have shown remarkable zero-shot inference in image classification, which naturally extends to semantic segmentation. They empower the segmentation pipeline to recognize seen and unseen categories, and users can segment almost any category in an image using prompt vocabulary [68, 81]. We divide current CLIP-based OVSS methods into two groups: training-required and training-free. The former allows models to be trained on some base classes in a supervised or weakly supervised manner. Typically, some works [16, 48, 52, 42] try to train a localization-aware CLIP which can naturally make dense predictions, while others [30, 13, 74, 73, 11, 38] select a subset of the CLIP’s pre-trained parameters and/or introduce a limited number of trainable parameters into the frozen CLIP, i.e., fine-tuning the CLIP to adapt to dense prediction on base classes. Still, training-free OVSS methods emphasize tapping into CLIP’s inherent localization capabilities with limited surgery of features or structures. MaskCLIP [79] pioneers the removal of query and key projections at the attention pooling layer of CLIP’s image encoder. Following it, subsequent studies [34, 63, 3, 28] adequately explore self-self attention (i.e., q-q, k-k or v-v self-attention), and these modifications somewhat mitigate noisy activations and spatial invariant perception of CLIP. Another stream [56, 23, 60, 2] is the two-stage method, which first generates category-agnostic mask proposals and then classifies the masks. Besides, some other foundation models (e.g. SAM [26], Stable Diffusion [54]) can be introduced to enhance the localization ability of CLIP, and these explorations also make sense [29, 2, 65].
Different from previous methods, we focus on the inherent characteristics of remote sensing images rather than the general attributes of natural images. The only contemporaneous work is [6], but it is training-required, like [74, 11]. Our SimFeatUp component, although it needs to be trained on a few images-only data beforehand, this process is independent of the semantic segmentation process and the trained weights can be used for almost any remote sensing data (like the foundation model in other works [29, 2]), so our method can still be seen as a training-free method.
@@ 3.1 CLIP
$\begin{array}[]{c}\boldsymbol{q}=\operatorname{Emb}_{q}(X),\boldsymbol{k}=\operatorname{Emb}_{k}(X),\boldsymbol{v}=\operatorname{Emb}_{v}(X),\\ \boldsymbol{y}=X+\operatorname{SA}\left(\boldsymbol{q},\boldsymbol{k},\boldsymbol{v}\right),\\ \boldsymbol{z}=\boldsymbol{y}+\operatorname{FFN}(\operatorname{LN}(\boldsymbol{y})),\end{array}$ (1)
@@ 4.1 SimFeatUp
FeatUp provides us with an excellent training paradigm for general upsamplers. However, it lacks some considerations for the training-free setting, leading it to be sub-optimal for the OVSS task, especially in remote sensing contexts.
@@ 4.2 Alleviating global bias
Figure 5: Comparison of before and after alleviating the global bias. (a) is the similarity map of patch tokens and cls tokens, some “non-building” regions also present high response, (b) is the original RGB image. Note that the right-hand histograms stretch the raw values for better presentation.
@@ 5.3 Comparison to State-of-the-art
Since the proposed SegEarth-OV is a training-free method and there is no previous OVSS method designed for remote sensing images, we select 5 state-of-the-art training-free OVSS models of natural images for comparison, including vanilla CLIP [51], MaskCLIP [79], SCLIP [63], GEM [3] and ClearCLIP [28].
Semantic segmentation. As listed in Tab. 1, SegEarth-OV achieves the best performance on all 8 semantic segmentation datasets. SegEarth-OV achieves more than 40% mIoU on 5 datasets and more than 50% on the UDD5 dataset, which implies that the OVSS method is feasible in remote sensing scenarios. Compared to the previous method, SegEarth-OV achieves a performance gain of more than 5% on 5 datasets and an average gain of 5.8% on 8 datasets. On the iSAID dataset, the mIoU of SegEarth-OV is only 21.7%, which is due to the fine-grained category delineation in this dataset, which covers 16 categories (see Appendix Tab. 7).
@@ 5.4 Ablation Study and Analysis
Remote sensing CLIPs for OVSS. We evaluate the performance of remote sensing CLIPs on OVSS, including RemoteCLIP [36], GeoRSCLIP [77], and SkyCLIP [66], which are trained on 0.8M, 5M, and 2.6M remote sensing data, respectively, without changing the model structure of CLIP. Since these works do not provide the ViT-B/16, we uniformly use ViT-B/32. Hence, we repeat the JBU operation 5 times in SimFeatUp. For fair comparison, we train the respective SimFeatUp for each model. As listed in Tab. 3, RemoteCLIP performs suboptimally to vanilla CLIP, which indicates that a small amount of domain data diminishes the model’s transfer capability. GeoRSCLIP achieves the best performance against SkyCLIP, which suggests that domain VLMs can benefit from more diverse domain-specific data. Moreover, the OVSS task effectively reflects the model’s discrimination and localization capabilities, and can serve as an evaluation metric for remote sensing VLMs.